• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kfuoti

    Log entry time 08:00:40 on September24,2005

    Entry number 155789

    keyword=Owl shift summary

    Owl Shift Summary:
    
    00:00 Shift begins while the first production run after Ryan and Lisa's tests 
    is underway.
    
    00:40 Profile scan started.  This is the first scan done after the tungsten 
    pre-radiator was removed.  I set the HV to -1600V, which is the PMTs' maximum.
    
    1:00  MCC says that they are still in the process of starting up Hall C again, 
    so wait to do bleedthrough measurement.
    
    1:50 Change IHWP to out
    
    6:45 IHWP in
    
    Cavity power for Compton is still seems low (270W).  
    
    
    
    
    Run Summary:
    
    IHWP IN
    ---------------------------------
    5318  Production;  42.7K
    5319  Profile Scan
    5320  Production;  36.9K
    
    
    IHWP OUT
    ---------------------------------
    5321  Production;  49.0K
    5322  Production;  34.8K
    5323  Production;  45.5K
    5324  Production;  46.1K
    5325  Production;  45.8K
    
    IHWP IN
    ----------------------------------
    5326  Production;   50K
    5327  Production....on to day shift